OpenStudy (anonymous):

Please help! secx=-2, find csc (x-pi/2)

OpenStudy (freckles):

sin is odd function so csc is odd function \[\csc(x-\frac{\pi}{2})=\csc(-(\frac{\pi}{2}-x))=-\csc(\frac{\pi}{2}-x)=-\sec(x) \text{ by co-function identity }\]

OpenStudy (solomonzelman):

you can start from new line in latex by pressing \\

OpenStudy (solomonzelman):

in boxes you need another trick, because in a box it won't work nicely (you can use a fraction in fraction even).